Solve each exponential equation. Give the exact value for x.
2^x=1/4
7^x=10
4^x=2.5
3^x=4

Answer by stanbon(75887)

Solve each exponential equation. Give the exact value for x.
2^x=1/4
2^x = 2^-2
x = -2
-----------------
7^x=10
x*log(7) = 1
x = 1/log(7)
x = 1.1833
-------------
4^x=2.5
xlog(4) = log(2.5)
x = log(2.5)/log(4) = 0.6610
--------------------
3^x=4
x = log(4)/log(3) = 1.2619
=================
Cheers,
Stan H.
=================